# Dust evolution during the protostellar collapse: influence on the coupling between the neutral gas and the magnetic field

### v1

@@ 2.2.2 Fragmentation and elastic properties
To establish a fragmentation condition, we follow Ormel et al. (2009) whose work involves dust grains in quiescent environments, that is to say porous aggregates colliding at subsonic velocities and made up of elementary, unbreakable pieces called monomers. Experimentally Blum & Wurm (2000) observed that complete fragmentation of the two colliding grains takes place as soon as the kinetic energy of the collision exceeds five times the total breaking energy $E_{\mathrm{br,tot}}=N_{\mathrm{tot}}E_{\mathrm{br}}$ of the grains (defined in Dominik & Tielens, 1997)
@@
where $N_{tot}\equiv\frac{m_{i}+m_{j}}{m_{\mathrm{mono}}}$ is the total number of electrostatic bonds between monomers. $E_{\mathrm{br}}$ is the energy required to break a single electrostatic bound and depends on the elastic properties of the grains, including the surface energy density $\gamma_{\mathrm{grain}}$ and the reduced elastic modulus $\varepsilon_{*}$. The latter is a function of the Poisson’s ratio and Young’s modulus, as detailed in Dominik & Tielens (1997). The mass of a monomer is related to the average monomer size as per $m_{\mathrm{mono}}\equiv\frac{4}{3}\pi s^{3}_{\mathrm{mono}}\rho_{\mathrm{grain}}$ where $\rho_{\mathrm{grain}}$ is the bulk intrinsic density of the grains. We use $A_{\mathrm{br}}=2.8\times 10^{3}$, which is experimentally measured in Blum & Wurm (2000). $s_{\mathrm{mono}}$ is the monomer size, taken to be equal to $100\ \mathrm{nm}$. Note that although $E_{\mathrm{br}}$ diminishes when reducing the monomer size, the total breaking energy actually increases due to the increasing number of bonds $N_{\mathrm{tot}}$. Consequently, reducing the monomer size leads to a higher fragmentation energy threshold.
@@ 3 Numerical methods
The numerical structure of the code used to treat the dust evolution has been thoroughly detailed in Lebreuilly et al. (2023). Here, we only shed light on the specific features relevant to our analysis. The version used here is the python sharkpy version. It accounts for dust evolution through coagulation and fragmentation processes, and computes the grain charge, ion and electron number density as well as the magnetic conductivities and resistivities of the mixture. In the scope of this work, the simulations were performed on a single cell, where physical properties can evolve with time, but with no spatial extension.
@@ 3.1.1 Coagulation and fragmentation
For each collision between grains i and j, we define a fragmentation coefficient $f_{\mathrm{frag},i,j}$ that is computed according to the fragmentation condition presented in Sect. 2.2.2 and as explained below. A fraction $(1-f_{\mathrm{frag},i,j})\,(m_{i}+m_{j})$ of the total mass involved in the collision is coagulated to form a larger grain. The rest of the mass $f_{\mathrm{frag},i,j}\,(m_{i}+m_{j})=m_{\mathrm{frag}}$ is shattered and redistributed as fragments among all the bins in the range $\left[m_{\mathrm{min}},0.1\ m_{\mathrm{frag}}\right]$, with $m_{\mathrm{min}}$ being the mass associated to the dust bin of size $s_{\mathrm{min}}$. The shattered mass $m_{\mathrm{frag}}$ is redistributed as a number density power-law: $\frac{dN}{ds}=s^{\psi}$, where $\psi=-3.5$.
@@ 3.1.2 Grain-grain erosion
Grain-grain erosion involves grains of very different sizes. Recently, Schrapler et al. (2018) conducted laboratory experiments to constrain the velocity threshold (for the onset of grain-grain erosion) and the corresponding efficiency by colliding micrometer projectiles onto a cm size target. They measured mass losses from the larger grain equal to a few times the mass of the smaller one in some cases. They observed the velocity threshold decreasing and the erosion efficiency rising when increasing the size difference between the projectile and the target. We assume their results can be translated to different grain size regimes (e.g., nanometer and micrometer grains at low density during the collapse). We implemented in our code the erosion velocity threshold (Fig. 8 in their paper) as a function of the grain size ratio $s_{\mathrm{ratio}}=s_{\mathrm{target}}/s_{\mathrm{projectile}}$ (where $s_{\mathrm{target}}$ is the size of the larger grain and $s_{\mathrm{projectile}}$ that of the smaller)
@@
and the erosion efficiency (Fig. 7 in their paper)
@@ 3.1.3 Electrostatic repulsion
Since grains are negatively charged on average, they experience electrostatic repulsion as they encounter, which modifies their effective collisional cross-section (see Eq. 13). In addition, we allow coagulation only if the kinetic energy of the colliding grains is higher than the Coulomb electrostatic energy ($E_{\mathrm{kinetic}}>E_{\mathrm{Coulomb}}$), which translates into a velocity threshold that needs to be exceeded for the colliding grains to stick together
@@ 3.1.4 Charging and resistivities
We use the scheme presented in Marchand et al. (2021) to compute the charge of each grain species and the number density of ions and electrons. It relies on different parameters such as the average ion ($\mathrm{HCO^{+}}$) mass $\mu_{\mathrm{ions}}=25$ , the size of ice coats $s_{\mathrm{ice}}=8\ \mathrm{nm}$, the sticking efficiency coefficient of electrons onto grains $s_{\mathrm{e}}=0.5$ and the cosmic ray ionization rate $\zeta=5\times 10^{-17}\mathrm{s^{-1}}$.
The conductivities (parallel, perpendicular and Hall) are inferred via the Hall factor (Eq. 2) and the individual conductivities of the different species
@@
To better understand the behavior of the resistivities, we include in the Appendix (Sect. A) the conductivity profiles, with an emphasis on the distinct influence of the different species, namely the electrons, the ions and the grains.
@@ 3.2.1 Magnetic field
We use an analytical expression that allows the magnetic field to increase with the density as per Li et al. (2011) and Marchand et al. (2016)
@@
where $B_{0}=30\ \mu\mathrm{G}$ is the magnetic field intensity at $n_{\mathrm{H}}=10^{4}\ \mathrm{cm^{-3}}$. We define the ambipolar diffusion timescale as
@@ 3.2.2 Collapse setup
The temperature is computed according to a barotropic equation used in Machida et al. (2006), Marchand et al. (2016) and Marchand et al. (2021)
@@
where the initial cloud temperature $T_{0}=10\ \mathrm{K}$ is a typical value measured in dense prestellar cores. $n_{\mathrm{H,ad}}$ is the threshold density that marks the transition between the isothermal and the adiabatic regime, and usually $n_{\mathrm{H,ad}}\simeq 10^{11}\ \mathrm{cm^{-3}}$. We consider a perfect gas with equation of state
@@
where $P_{\mathrm{g}}$ is the gas pressure, $\rho$ the mass density of the gas, $\mu_{\mathrm{g}}=2.3$ the mean molecular weight and $m_{\mathrm{H}}$ the mass of a hydrogen atom. The sound speed is given by $c_{\mathrm{s},0}\equiv\sqrt{\frac{\gamma P_{\mathrm{g}}}{\rho}}=\sqrt{\frac{\gamma k_{\mathrm{B}}T_{0}}{\mu_{\rm{g}}m_{\rm{H}}}}$ where the effective adiabatic index is taken to be that of a monoatomic gas $\gamma=\frac{5}{3}$ since the temperature of $H_{2}$ molecules is not high enough to excite vibrational and rotational energy states in this cold environment. Figure 1 gives the gas density evolution with time during the protostellar collapse. The data is extracted from the central cell of a shark 1D collapse simulation without rotation or magnetic support. The collapse is initiated and halted exactly as in Lebreuilly et al. (2023), that is to say with an initial density of $\rho=9.2\times 10^{-18}\ \mathrm{g\ cm^{-3}}\left(n_{\mathrm{H}}=2.4\times 10^{6}\ \mathrm{cm^{-3}}\right)$ and a final density of $\rho=10^{-10}\ \mathrm{g\,cm^{-3}}\left(n_{\mathrm{H}}=2.6\times 10^{13}\ \mathrm{cm^{-3}}\right)$, which corresponds to a solar mass cloud associated to a gravitational to thermal energy ratio $\alpha=0.25$.
We supplement in the Appendix (Sect. 12) the comparison between a non-evolving MRN dust distribution, a growing distribution and a growing plus fragmenting distribution at different instants of the collapse. We recover the results of Lebreuilly et al. (2023) and Kawasaki et al. (2022).
@@ 3.2.3 Parameters investigated
Table 1 displays the parameters explored and their corresponding values. We highlight in blue in the first column the reference values used when a given parameter is not being varied (note that we use two reference values for the grain surface energy $\gamma_{\mathrm{grain}}$). From those reference values, we vary the parameters independently, one at a time.
  1. 1.
    $\gamma_{\mathrm{grain}}$ is the surface energy of the grains, which controls the breaking energy. Given the uncertainties around the elastic properties and tensile strength of the grains, and given that there is a temperature dependency of the surface energy of the grains, we vary the latter between two extreme values measured in Musiolik & Wurm (2019): $\gamma_{\mathrm{grain}}=2.9\ \mathrm{erg\ cm^{-2}}$ and $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$. Using the elastic modulus for icy-grains to be $\varepsilon_{*}=3.7\times 10^{10}$, the two extreme values for the surface energy correspond respectively to the case of minimum breaking energy ($E_{\mathrm{{br,min}}}=1.27\times 10^{-10}\ \mathrm{erg}$) and maximum breaking energy ($E_{\mathrm{{br,max}}}=4.11\times 10^{-7}\ \mathrm{erg}$) we could encounter.
  2. 2.
    erosion is a boolean that activates or deactivates the grain-grain erosion process.
  3. 3.
    $\mathrm{repulsion}$ is a boolean that activates or deactivates the grain-grain electrostatic repulsion process.
  4. 4.
    $\delta_{\mathrm{AD}}$ is the ambipolar drift intensity, taking values between 0.1 and 10. Those limit values are expected in collapsing environments, as suggested by Fig. D.
  5. 5.
    $B_{0}$ is the initial magnetic field strength corresponding to a gas density of $n_{\mathrm{H}}=10^{4}\ \mathrm{cm^{-3}}$ in the scaling approximation of Eq. (21). It takes values between $10^{-5}\ \mathrm{G}$ and $5\times 10^{-5}\ \mathrm{G}$, which are common values measured in dense cores, see for example Troland & Crutcher (2008).
  6. 6.
    $\zeta$ is the ionization rate induced by cosmic rays. (see Sect. 3.1.4). It takes values between $5\times 10^{-18}\ \mathrm{s^{-1}}$ and $5\times 10^{-16}\ \mathrm{s^{-1}}$, which are typical values measured in Padovani et al. (2022). Constraining this parameter is a daunting task due to the uncertainties in the chemical networks used and to observation difficulties.
  7. 7.
    $\epsilon_{\mathrm{stick}}$ is a grain sticking coefficient which affects the sticking probability between two colliding grains, but not that of fragmentation. It takes values between $0.1$ and $1$. Since this parameter is poorly constrained and dependent on the grain surface energy and grain coats nature, we explored a wide range of values.
@@ 4 Results
Figure 2: Dust distributions at different stages during the collapse. First row: mass dust-to-gas ratio. Second row: number dust-to-gas ratio. Third row: resistivities profile and ion/electron numerical densities (Hall resistivity: dotted line for the negative values and solid line for the positive ones). Dashed black lines: reference resistivities obtained with a fixed MRN dust distribution. Fourth row: parallel, perpendicular and Hall conductivities and evolution of the peak size of the dust distribution. Comparison between different values of the grain surface energy expressed in cgs units ($\mathrm{erg\ cm^{-2}}$).
@@
Figure 3: Likewise Fig. 2 but investigating the influence of grain-grain erosion with $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$.
@@
Figure 4: ambipolar resistivity $\eta_{\mathrm{AD}}$ evolution during the protostellar collapse for different values of the ambipolar drift intensity or equivalently different ambipolar drift velocities: $\delta_{\mathrm{AD}}=0.1\ (0.1V_{\mathrm{AD}})$, $\delta_{\mathrm{AD}}=1\ (V_{\mathrm{AD}})$, and $\delta_{\mathrm{AD}}=10\ (10V_{\mathrm{AD}})$. Grain-grain electrostatic repulsion (first column) and erosion (second column) are also included (dotted lines). Second row: likewise but with a sticking efficiency of 0.5.
Parameters exploration
$\gamma_{\mathrm{grain}}(\mathrm{erg\ cm^{-2}})$ $[2.9,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}10},30,50,100,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}190}]$
$\mathrm{erosion}$ $[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{False}},\mathrm{True}]$
$\mathrm{repulsion}$ $[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{False}},\mathrm{True}]$
$\delta_{\mathrm{AD}}$ $[0.1,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1},10]$
$B_{0}(\mathrm{\mu G})$ $[10,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}30},50]$
$\zeta(\mathrm{s^{-1}})$ $[5\times 10^{-18},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5\times 10^{-17}},5\times 10^{-16}]$
$\epsilon_{\mathrm{stick}}$ $[0.1,0.5,0.7,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1}]$
@@
Table 1: Range of values taken by the explored parameters. In blue are the reference values used when a given parameter is not being varied. Two reference values have been used for the grain surface energy, corresponding to fragile grains on one hand, and to resilient grains on the other.
@@
We first explored the impact of the surface energy of the grains. Then, we found that the degree of influence of the other parameters depends strongly on the fragmentation rate, that is on the grain surface energy. Consequently, we do not make any assumption regarding the composition of the grains but rather choose to investigate the influence of the other parameters for two different extreme values of the grain surface energy, corresponding to resistant grains ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$) on one hand, and fragile ones ($\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$) on the other hand. We investigate the influence of the ambipolar drift intensity and its combined effect with electrostatic repulsion and grain-grain erosion for a grain surface energy of $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$. The analysis of the grain-grain sticking efficiency $\epsilon_{\mathrm{stick}}$ is attached to the Appendix (Sect. 17) and was carried out with $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$. The initial magnetic field strength $B_{0}$ and the cosmic-ray ionization rate $\zeta$ are mentioned and commented on in Sect. 5 ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$ if not specified otherwise).

### v2

@@ 2.2.2 Fragmentation and elastic properties
To establish a fragmentation condition, we follow Ormel et al. (2009), whose work involves dust grains in quiescent environments, that is to say: porous aggregates colliding at subsonic velocities and made up of elementary, unbreakable pieces called monomers. Experimentally Blum & Wurm (2000) observed that complete fragmentation of the two colliding grains takes place as soon as the kinetic energy of the collision exceeds five times the total breaking energy, $E_{\mathrm{br,tot}}=N_{\mathrm{tot}}E_{\mathrm{br}}$, of the grains (defined in Dominik & Tielens, 1997):
@@
where $N_{tot}\equiv\frac{m_{i}+m_{j}}{m_{\mathrm{mono}}}$ is the total number of electrostatic bonds between monomers. $E_{\mathrm{br}}$ is the energy required to break a single electrostatic bound and depends on the elastic properties of the grains, including the surface energy density $\gamma_{\mathrm{grain}}$ and the reduced elastic modulus $\varepsilon_{*}$. The latter is a function of the Poisson’s ratio and Young’s modulus, as detailed in Dominik & Tielens (1997). The mass of a monomer is related to the average monomer size as per $m_{\mathrm{mono}}\equiv\frac{4}{3}\pi s^{3}_{\mathrm{mono}}\rho_{\mathrm{grain}}$ where $\rho_{\mathrm{grain}}$ is the bulk intrinsic density of the grains. We use $A_{\mathrm{br}}=2.8\times 10^{3}$, which is experimentally measured in Blum & Wurm (2000). Here, $s_{\mathrm{mono}}$ is the monomer size, taken to be equal to $100\ \mathrm{nm}$. We note that although $E_{\mathrm{br}}$ diminishes when reducing the monomer size, the total breaking energy actually increases due to the increasing number of bonds $N_{\mathrm{tot}}$. Consequently, reducing the monomer size leads to a higher fragmentation energy threshold.
@@ 3 Numerical methods
The numerical structure of the code used to treat the dust evolution has been thoroughly detailed in Lebreuilly et al. (2023). Here, we only shed light on the specific features relevant to our analysis. The version used here is the python sharkpy version. It accounts for dust evolution through coagulation and fragmentation processes, and computes the grain charge, ion, and electron number density as well as the magnetic conductivities and resistivities of the mixture. Within the scope of this work, the simulations were performed on a single cell, where physical properties can evolve with time, but with no spatial extension.
@@ 3.1.1 Coagulation and fragmentation
For each collision between grains i and j, we define a fragmentation coefficient $f_{\mathrm{frag},i,j}$ that is computed according to the fragmentation condition presented in Sect. 2.2.2 (and explained below). A fraction $(1-f_{\mathrm{frag},i,j})\,(m_{i}+m_{j})$ of the total mass involved in the collision is coagulated to form a larger grain. The rest of the mass, $f_{\mathrm{frag},i,j}\,(m_{i}+m_{j})=m_{\mathrm{frag}}$, is shattered and redistributed as fragments among all the bins in the range $\left[m_{\mathrm{min}},0.1\ m_{\mathrm{frag}}\right]$, with $m_{\mathrm{min}}$ being the mass associated with the dust bin of size $s_{\mathrm{min}}$. The shattered mass $m_{\mathrm{frag}}$ is redistributed as a number density power-law: $\frac{dN}{ds}=s^{\psi}$, where $\psi=-3.5$.
@@ 3.1.2 Grain-grain erosion
Grain-grain erosion involves grains of very different sizes. Recently, Schrapler et al. (2018) conducted laboratory experiments to constrain the velocity threshold (for the onset of grain-grain erosion) and the corresponding efficiency by colliding micrometer projectiles onto a cm size target. They measured mass losses from the larger grain equal to a few times the mass of the smaller one in some cases. They observed the velocity threshold decreasing and the erosion efficiency rising when increasing the size difference between the projectile and the target. We assume their results can be translated to different grain size regimes (e.g. nanometer and micrometer grains at low density during the collapse). We implemented in our code the erosion velocity threshold (Fig. 8 in their paper) as a function of the grain size ratio $s_{\mathrm{ratio}}=s_{\mathrm{target}}/s_{\mathrm{projectile}}$ (where $s_{\mathrm{target}}$ is the size of the larger grain and $s_{\mathrm{projectile}}$ that of the smaller)
@@
and the erosion efficiency (Fig. 7 in their paper):
@@ 3.1.3 Electrostatic repulsion
Since grains are negatively charged on average, they experience electrostatic repulsion as they encounter, which modifies their effective collisional cross-section (see Eq. 13). In addition, we allow for coagulation only if the kinetic energy of the colliding grains is higher than the Coulomb electrostatic energy ($E_{\mathrm{kinetic}}>E_{\mathrm{Coulomb}}$). This translates into a velocity threshold that needs to be exceeded for the colliding grains to stick together:
@@ 3.1.4 Charging and resistivities
We used the scheme presented in Marchand et al. (2021) to compute the charge of each grain species and the number density of ions and electrons. It relies on different parameters such as the average ion ($\mathrm{HCO^{+}}$) mass $\mu_{\mathrm{ions}}=25$ , the size of ice coats $s_{\mathrm{ice}}=8\ \mathrm{nm}$, the sticking efficiency coefficient of electrons onto grains $s_{\mathrm{e}}=0.5$ and the cosmic-ray ionisation rate $\zeta=5\times 10^{-17}\mathrm{s^{-1}}$.
The conductivities (parallel, perpendicular and Hall) are inferred via the Hall factor (Eq. 2) and the individual conductivities of the different species:
@@
To better understand the behavior of the resistivities, in Appendix A, we include the conductivity profiles, with an emphasis on the distinct influence of the different species, namely, the electrons, the ions and the grains.
@@ 3.2.1 Magnetic field
We use an analytical expression that allows for the magnetic field to increase with the density as per Li et al. (2011) and Marchand et al. (2016):
@@
where $B_{0}=30\ \mu\mathrm{G}$ is the magnetic field intensity at $n_{\mathrm{H}}=10^{4}\ \mathrm{cm^{-3}}$. We define the ambipolar diffusion timescale as:
@@ 3.2.2 Collapse setup
The temperature is computed according to a barotropic equation used in Machida et al. (2006), Marchand et al. (2016) and Marchand et al. (2021):
@@
where the initial cloud temperature $T_{0}=10\ \mathrm{K}$ is a typical value measured in dense prestellar cores. Also, $n_{\mathrm{H,ad}}$ is the threshold density that marks the transition between the isothermal and the adiabatic regime, usually $n_{\mathrm{H,ad}}\simeq 10^{11}\ \mathrm{cm^{-3}}$. We consider a perfect gas with equation of state:
@@
where $P_{\mathrm{g}}$ is the gas pressure, $\rho$ is the mass density of the gas, $\mu_{\mathrm{g}}=2.3$ is the mean molecular weight, and $m_{\mathrm{H}}$ is the mass of a hydrogen atom. The sound speed is given by $c_{\mathrm{s},0}\equiv\sqrt{\frac{\gamma P_{\mathrm{g}}}{\rho}}=\sqrt{\frac{\gamma k_{\mathrm{B}}T_{0}}{\mu_{\rm{g}}m_{\rm{H}}}}$, where the effective adiabatic index is taken to be that of a monoatomic gas $\gamma=\frac{5}{3}$ since the temperature of $H_{2}$ molecules is not high enough to excite vibrational and rotational energy states in this cold environment. Figure 1 gives the gas density evolution with time during the protostellar collapse. The data is extracted from the central cell of a shark 1D collapse simulation without rotation or magnetic support. The collapse is initiated and halted exactly as in Lebreuilly et al. (2023), that is, to say with an initial density of $\rho=9.2\times 10^{-18}\ \mathrm{g\ cm^{-3}}\left(n_{\mathrm{H}}=2.4\times 10^{6}\ \mathrm{cm^{-3}}\right)$ and a final density of $\rho=10^{-10}\ \mathrm{g\,cm^{-3}}\left(n_{\mathrm{H}}=2.6\times 10^{13}\ \mathrm{cm^{-3}}\right)$, which corresponds to a solar mass cloud associated with a gravitational to thermal energy ratio $\alpha=0.25$.
In Appendix 12, we offer a comparison between a non-evolving MRN dust distribution, a growing distribution and a growing plus fragmenting distribution at different instants of the collapse. We recovered the results of Lebreuilly et al. (2023) and Kawasaki et al. (2022).
@@ 3.2.3 Parameters investigated
Table 1 displays the parameters explored and their corresponding values. We highlight in bold in the first column the reference values used when a given parameter is not being varied (note: we used two reference values for the grain surface energy $\gamma_{\mathrm{grain}}$). From those reference values, we vary the parameters independently, one at a time.
1. $\gamma_{\mathrm{grain}}$ is the surface energy of the grains, which controls the breaking energy. Given the uncertainties around the elastic properties and tensile strength of the grains, and given that there is a temperature dependency of the surface energy of the grains, we vary the latter between two extreme values measured in Musiolik & Wurm (2019): $\gamma_{\mathrm{grain}}=2.9\ \mathrm{erg\ cm^{-2}}$ and $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$. Using the elastic modulus for icy-grains to be $\varepsilon_{*}=3.7\times 10^{10}$, the two extreme values for the surface energy correspond respectively to the case of minimum breaking energy ($E_{\mathrm{{br,min}}}=1.27\times 10^{-10}\ \mathrm{erg}$) and maximum breaking energy ($E_{\mathrm{{br,max}}}=4.11\times 10^{-7}\ \mathrm{erg}$) we could encounter.
2. Erosion is a boolean that activates or deactivates the grain-grain erosion process.
3. $\mathrm{Repulsion}$ is a boolean that activates or deactivates the grain-grain electrostatic repulsion process.
4. $\delta_{\mathrm{AD}}$ is the ambipolar drift intensity, taking values between 0.1 and 10. Those limit values are expected in collapsing environments, as suggested by Fig. D.
5. $B_{0}$ is the initial magnetic field strength corresponding to a gas density of $n_{\mathrm{H}}=10^{4}\ \mathrm{cm^{-3}}$ in the scaling approximation of Eq. (21). It takes values between $10^{-5}\ \mathrm{G}$ and $5\times 10^{-5}\ \mathrm{G}$, which are common values measured in dense cores, see for example Troland & Crutcher (2008).
6. $\zeta$ is the ionisation rate induced by cosmic rays. (see Sect. 3.1.4). It takes values between $5\times 10^{-18}\ \mathrm{s^{-1}}$ and $5\times 10^{-16}\ \mathrm{s^{-1}}$, which are typical values measured in Padovani et al. (2022). Constraining this parameter is a daunting task due to the uncertainties in the chemical networks used and to observation difficulties.
7. $\epsilon_{\mathrm{stick}}$ is a grain sticking coefficient which affects the sticking probability between two colliding grains, but not that of fragmentation. It takes values between $0.1$ and $1$. Since this parameter is poorly constrained and dependent on the grain surface energy and grain coats nature, we explored a wide range of values.
@@ 4 Results
Figure 2: Dust distributions at different stages during the collapse. First row: Mass dust-to-gas ratio. Second row: Number dust-to-gas ratio. Third row: Resistivities profile and ion and electron numerical densities (Hall resistivity: dotted line for the negative values and solid line for the positive ones). Dashed black lines: Reference resistivities obtained with a fixed MRN dust distribution. Fourth row: Parallel, perpendicular, and Hall conductivities, and the evolution of the peak size of the dust distribution. Comparison between different values of the grain surface energy expressed in cgs units ($\mathrm{erg\ cm^{-2}}$).
@@
Figure 3: Same as Fig. 2, but investigating the influence of grain-grain erosion with $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$.
@@
Figure 4: Ambipolar resistivity $\eta_{\mathrm{AD}}$ evolution during the protostellar collapse for different values of the ambipolar drift intensity or equivalently different ambipolar drift velocities: $\delta_{\mathrm{AD}}=0.1\ (0.1V_{\mathrm{AD}})$, $\delta_{\mathrm{AD}}=1\ (V_{\mathrm{AD}})$, and $\delta_{\mathrm{AD}}=10\ (10V_{\mathrm{AD}})$. Grain-grain electrostatic repulsion (first column) and erosion (second column) are also included (dotted lines). Second row: Same details, but with a sticking efficiency of 0.5.
Table 1: Range of values taken by the explored parameters.
Parameter exploration
$\gamma_{\mathrm{grain}}(\mathrm{erg\ cm^{-2}})$ $[2.9,\boldsymbol{10},30,50,100,\boldsymbol{190}]$
$\mathrm{erosion}$ $[\boldsymbol{\mathrm{False}},\mathrm{True}]$
$\mathrm{repulsion}$ $[\boldsymbol{\mathrm{False}},\mathrm{True}]$
$\delta_{\mathrm{AD}}$ $[0.1,\boldsymbol{1},10]$
$B_{0}(\mathrm{\mu G})$ $[10,\boldsymbol{30},50]$
$\zeta(\mathrm{s^{-1}})$ $[5\times 10^{-18},\boldsymbol{5\times 10^{-17}},5\times 10^{-16}]$
$\epsilon_{\mathrm{stick}}$ $[0.1,0.5,0.7,\boldsymbol{1}]$
@@
 [^1] In bold are the reference values used when a given parameter is not being varied. Two reference values have been used for the grain surface energy, corresponding to fragile grains on one hand, and to resilient grains on the other.
@@
We first explored the impact of the surface energy of the grains. Then, we found that the degree of influence of the other parameters depends strongly on the fragmentation rate, that is on the grain surface energy. Consequently, we do not make any assumption regarding the composition of the grains but rather choose to investigate the influence of the other parameters for two different extreme values of the grain surface energy, corresponding to resistant grains ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$) on one hand, and fragile ones ($\gamma_{\mathrm{grain}}=10\ \mathrm{erg\ cm^{-2}}$) on the other hand. We investigate the influence of the ambipolar drift intensity and its combined effect with electrostatic repulsion and grain-grain erosion for a grain surface energy of $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$. The analysis of the grain-grain sticking efficiency $\epsilon_{\mathrm{stick}}$ is given in the appendix (Appendix 17), performed with $\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$. The initial magnetic field strength $B_{0}$ and the cosmic-ray ionisation rate $\zeta$ are mentioned and commented on in Sect. 5 ($\gamma_{\mathrm{grain}}=190\ \mathrm{erg\ cm^{-2}}$ if not specified otherwise).